I need help with theese two question, Please.
1. The Nutty Professor store wishes to mix peanuts that sell for $2 per pound with cashews that sell for $6 per pound. The store plans to use five pounds more of peanuts than cashews and sell the mixture for $26. How many pounds of each nut should be used in the mixture?
AND this question too!
2. A stock portfolio is currently valued at $5872.44. IBM is currently selling at $82.01 per share and Microsoft is currently selling at $26.52 per share. If the number of IBM shares is 25 less than the number of Microsoft shares, find the number of shares of each.

Answer by stanbon(48564)

1. The Nutty Professor store wishes to mix peanuts that sell for $2 per pound with cashews that sell for $6 per pound. The store plans to use five pounds more of peanuts than cashews and sell the mixture for $26. How many pounds of each nut should be used in the mixture?
-----------------
Let amount of cashews be "x" lbs ; Value of these is 6x dollars
Amount of peanuts is "x+5" lbs ; Value of these is 2(x+5)=2x+10 dollars
EQUATION:
value + value = 26 dollars
6x+2x+10=26
8x=16
x=2 lbs. (amount of cashews in the mixture)
x+5=7 lbs.(amount of cashews in the mixture)
------------------------
2. A stock portfolio is currently valued at $5872.44. IBM is currently selling at $82.01 per share and Microsoft is currently selling at $26.52 per share. If the number of IBM shares is 25 less than the number of Microsoft shares, find the number of shares of each.
-------
Let number of Micro shares be "x" ; Value of these is 26.52x dollars
Number of IBM shares is "x-25" ; Val of these is 82.01(x-25)=$(82.01x-2050.25)
EQUATION:
value + value=5872.44 dollars
26.52x+82.01x-2050.25=5872.44
108.53x=7922.69
x=73 (# of shares of Microsoft)
x-25=48 (# of shares of IBM)
Cheers,
Stan H.